OpenStudy (anonymous):

SAT problem: For what values of a and b are the following equations both true? a+2b+1 -3a-8b=1 I got to where i have gotten only one variable and i am looking for a. Please explain in detail!

OpenStudy (anonymous):

*a+2b=1

OpenStudy (anonymous):

a=3/7, b=-4/7

OpenStudy (anonymous):

sorry wrong

OpenStudy (anonymous):

the book says after eliminating one of the variables you get -3a-1=4-4a but i have no clue where THAT equation came from

OpenStudy (anonymous):

2b=1-a. so 8b=4(1-a)

OpenStudy (anonymous):

sub 8b in eq2

myininaya (myininaya):

multiply the first equation by 3 to receive 3a+6b=3 now add to the other equation: +(-3a-8b=1) -------------- 0a-2b=4 -2b=4 b=-2

OpenStudy (anonymous):

to eliminate one of them, try multiplying the first equation by 4 4a + 8b = 4 -3a -8b = 1 now add both equations together a = 5 b = -2

myininaya (myininaya):

now plug b=-2 into one of equations given to you and solve for a

OpenStudy (anonymous):

web site is taking a break. me too!